OpenStudy (gabylovesyou):

Find the non-extraneous solutions of the square root of the quantity (x plus 9) minus 5 equals quantity x plus 4. x = −2 x = −9 and x = −2 x = 9 x = −9 and x = −8

OpenStudy (gabylovesyou):

\[\sqrt{x + 9}- 5 = x + 4\]

OpenStudy (gabylovesyou):

@eta

OpenStudy (anonymous):

Add \(5\) to both the sides..

OpenStudy (amorfide):

you want to isolate the root so that you can square both sides which you can then form a quadratic and solve for x substitute your x values into the original equation and you will find that only one of them or both of them will work

OpenStudy (anonymous):

And after that, square both the sides..

OpenStudy (gabylovesyou):

x = -9 ?

OpenStudy (anonymous):

Oh, let me check, I have not solved it..

OpenStudy (gabylovesyou):

wait i got x^2 + 18x + 81

OpenStudy (amorfide):

for the right hand side yes

OpenStudy (anonymous):

\[x+9 = (x+9)^2\]

OpenStudy (amorfide):

subtract the left hand side from the right hand side to make it equal to 0

OpenStudy (amorfide):

then solve the quadratic

OpenStudy (anonymous):

Do you have answer?

OpenStudy (anonymous):

Have not you got this?